# A Moment-Conserving Discontinuous Galerkin Representation of the Relativistic Maxwellian Distribution

### v1

@@ 3.2 Robust Algorithm for Computing DG Moments
Numerically, strong gradients in the drift velocity may be super-luminal even if the cell average is less than the speed of light. Constraining the drift velocity to be well-behaved is even harder when trying to compute $|\mathbf{v}_{b}|^{2}$ in the Lorentz boost factor and the rest-frame density. To solve this issue, we applied the following algorithm to the relativistic solver to utilize the spatial component of the bulk four-velocity in the moment calculation. The spatial component of the bulk four-velocity is guaranteed to be well behaved as it does not have to be bounded from above, and computing the needed Lorentz boost factor from this quantity is also guaranteed to be positive definite. To achieve this refactor the steps are now:
@@
    Compute $\mathbf{v}_{b}$ with equation 11 using weak division.
@@
      Evaluate $\mathbf{v}_{b}$ at polynomial order 1 Gauss-Lobatto points (the cell vertices).
@@
    Compute the rest frame density, $n$, using weak multiplication: $n=\frac{N}{\Gamma}$.
@@
    $P=nT=\frac{m_{0}}{d}\int f^{\prime}(\mathbf{u^{\prime}})\left(\Gamma_{\mathbf{u}}^{2}\gamma^{\prime}-2\Gamma_{\mathbf{u}}\mathbf{u}_{b}\cdot\mathbf{u}^{\prime}+\frac{(\mathbf{u_{b}}\cdot\mathbf{u^{\prime}})^{2}-1}{\gamma^{\prime}}\right)d^{3}u^{\prime}.$ (23)
@@ 3.3 Maxwell-Jüttner Initialization Routine
Initialization of the MJ distribution begins with the spatially DG-represented quantities $n,\mathbf{v}_{b}$, $T$ and $\Gamma$ computed via the routine in Section 3.2. These quantities are then used to calculate the MJ distribution, equation 2, at the quadrature points. We can then transform the nodal representation onto the modal phase basis, $\phi(\mathbf{x},\mathbf{u})$.
@@
However, directly evaluating the MJ distribution, equation 2, onto the momentum grid has two complications. First, only a limited range of temperatures can be computable with finite precision because of the exponential behavior of the modified Bessel function. It is thus impossible to initialize the distribution for non-relativistic and ultra-relativistic temperatures. Second, in terms of dimension, $d$, the modified Bessel function takes the form $K_{\frac{d+1}{2}}$, which is (potentially) of fractional order. These are notoriously difficult to compute accurately to the desired finite precision.
@@
To avoid the issues of calculating the modified Bessel functions altogether, we replace them with their asymptotic expansions. To leading order, the expansions of $K_{1},K_{3/2},$ and $K_{2}$ are all identical, $K_{\nu}(x)\sim\sqrt{\pi/2x}\exp{(-x)}$ where $\nu=(1,3/2,2)$. They each retain the same leading-order asymptotic behavior as $x\to 0,K_{\nu}(x)\to\infty$ and as $x\to\infty$, $K_{\nu}(x)\to 0$. This means that the projected distribution function is offset in magnitude (the density) only. This unnormalized distribution, expanding $K_{\nu}$ in the Maxwell-Jüttner distribution, equation 2, is written as
@@
$\widetilde{f}^{\prime MJ}(\mathbf{u}^{\prime})=\frac{n}{2c(2\pi)^{\frac{d-1}{2}}(m_{0}c)^{d}\theta^{\frac{d-1}{2}}\sqrt{\pi\theta/2}}\exp\left(\frac{1}{\theta}-\left(\frac{\Gamma(\gamma(\mathbf{p^{\prime}})-\mathbf{v}_{b}\cdot\mathbf{p}^{\prime}/(m_{0}c^{2}))}{\theta}\right)\right).$ (24)
We retain the normalization term in practice because it improves convergence of higher moments. However, it may also be absorbed into the density normalization constant: $\widetilde{f}^{\prime MJ}(\mathbf{u}^{\prime})=n\exp(...)$. Which is useful when the normalization is too cumbersome to compute or unknown.
@@
The density moment of equation 24 can be corrected by multiplying the ratio computed via weak division of the correct density moment over the asymptotic density moment to give the properly normalized distribution. Correcting with this density ratio allows us to avoid finite precision errors associated with the modified Bessel functions and simultaneously corrects any errors in the density introduced by the discrete projection.
@@
There are two common cases we consider for this density correction. The first case arises during initial conditions, when we know the input stationary frame density and thus can simply use that desired density moment to correct the distribution function and avoid errors in both the modified Bessel functions and projections. The second case is for operators such as the BGK collision operator, where the density is a time-dependent quantity. In this case, we compute the moments described in equations 10 and 11 to obtain the stationary frame density of the evolving distribution function, and then use that stationary frame density in the correction routine to ensure that the projected MJ distribution has the exact density of the evolving distribution.
@@
In discrete, finite grids, the higher moments of the MJ projection, $\mathbf{v}_{b}$ and $T$, may also deviate significantly from the desired values. To reduce the error between the discretely projected distributions moments, we employ an iterative method. Starting with the desired moments, $n_{0}$, $\mathbf{v}_{b,0}$, and $T_{0}$ we create a discrete projection $\tilde{f}$. The moments of this discrete distribution $\tilde{n}$, $\tilde{\mathbf{v}}_{b}$, and $\tilde{T}$, will not necessarily match $n_{0}$, $\mathbf{v}_{b,0}$, and $T_{0}$. We then use Picard iteration to minimize the difference between these sets of moments. For an iteration $k$, and a vector of the moments $\mathbf{M}=\{n,\mathbf{v}_{b},T\}$ we wish to converge, we have an iterations scheme
@@
An alternative algorithm, presented in Section 3.3 of Dzanic et al. (2023), employs a Newton method which, in some cases, would reduce the number of required iterations. This algorithm is extended to the relativistic Maxwell-Jüttner by replacing the distribution function, using the moments outlined in Section 2.1, and recomputing the partials needed by the Jacobian. However, we found that, for modal DG schemes, the Newton method is prohibitively expensive.  [^3] For the non-relativistic BGK operator, Dzanic, Witherden and Martinelli Dzanic et al. (2023) report the Newton correction typically takes 1-2 iterations to reach machine precision. Such an approach could also apply to the Maxwell-Jüttner distribution. However, for specifically the modal DG represented quantities, the Jacobian required by this method would be prohibitively expensive to compute, and each cell would need to invert the Jacobian of the size number of basis, times the number of moments, squared. The trade off is that the corrective method provided here will take additional iterations but is simpler to implement and can automatically handle arbitrary equilibrium functions, such as MJ, Maximilian’s, and curved spaces.
@@ 4 Tests
This section explores the accuracy of projecting the MJ distribution function and tests the combination of the projection routine and the moment routine. We address the accuracy of projecting a MJ distribution, examining grid requirements for capturing the distribution within a certain tolerance of error. We demonstrate how the correction routine alters the discrete distribution to match the moments. We also provide analytic estimates for maximum and minimum temperatures representable by the finite momentum grid. Finally, we integrate these concepts into a conservative relativistic BGK operator.
@@ 4.1 Convergence Test
Figure 1: Absolute error between the projected Maxwell-Jüttner distribution’s moments and the desired moments. The moment values here are $n=1.0,v_{b}=0.5$, and $T=1.0$ and the momentum bounds extend from $u_{max}=\pm 160$.
@@
From Figure 1, we see that convergence to machine precision of the beam velocity and temperature takes around 10,000 momentum cells with moderate temperature and beam velocity of 1.0 and 0.5, respectively. Convergence indicates we have sufficiently wide momentum bounds to resolve the distribution function for the specified grid size and moments. However, such a large number of grid points is impractical for an actual simulation, hence the need for a moment-correction routine.
The requirement for having a large momentum domain of $u_{max}=\pm 160$ in Figure 1 arises from the difference in the MJ distribution’s asymptotic behavior for large values of momentum, $u$. Compared with the Maxwellian distribution, this difference becomes apparent. For a 1D Maxwellian distribution, $f^{Max}\sim\exp(-mv^{2}/kT)$. In contrast, the 1D Maxwell-Jüttner goes as $f^{MJ}\sim\exp(-m_{0}c^{2}|u|/kT)$ when $m_{0}c^{2}|u|/kT\gg 1$. Hence, the Maxwellian approaches zero much faster than the Maxwell-Jüttner on their respective grids. This detail leads to some subtle issues if not minded. To illustrate, the bounds of $u_{max}=\pm 10$ may seem reasonable. However, we are truncating more distribution function than is visually apparent. For example, the lost density fraction for the bounds $u_{max}=\pm 10$ is ($\delta N/N$) $\sim 0.003685$.
@@
Figure 2: MJ plotted with the error between the corrected Maxwell-Jüttner distribution $f_{c}$ and the uncorrected distribution $f_{unc}$. Since the corrected and uncorrected are indistinguishable on the plot, only the corrected distribution is plotted here.
@@ 4.2 Representation Limits of Moments of the MJ Distribution
Figure 3 panel (a) scans the case where the bulk of the distribution is non-relativistic for a 1D momentum space. Contours of the number of iterations required to correct the moments in $\Delta u$ vs temperature $T$ demonstrate that the convergence of temperature is not guaranteed for $T<T_{min,non-rel.}$. Where
$\frac{T_{min,non-rel.}}{m_{0}c^{2}}=\frac{\Delta u^{2}}{2(p+1)^{2}}$ (26)
and $\Delta u$ is the grid spacing in momentum space and $p$ is the poly-order of the DG scheme. Scanning from warmer to cooler temperatures, the number of iterations required to converge the scheme increases. Eventually, the algorithm hits the maximum number of iterations, and occasionally fails to converge below $T_{min,non-rel.}$. The grid spacing and number of iterations effectively set the minimum grid temperature for the MJ. The trend of minimum temperature holds across 1D, 2D, and 3D momentum spaces.
@@
$\frac{T_{max}}{m_{0}c^{2}}=\frac{u_{max}\sqrt{1+u_{max}^{2}}-\text{asinh}(u_{max})}{2u_{max}}$ (27)
by computing the temperature of a flat distribution between finite bounds of equation 13 where we have assumed the distribution has no net flow.
In the relativistic limit, Figure 3 panel (b), we can simultaneously show the maximum MJ temperature $T_{max,rel.}$ and the minimum temperature, $T_{min,rel.}$. Scanning $u_{max}$ and $T$, a clear region appears where the temperature is able to converge. Like in the non-relativistic case, the maximum temperature projectable by a MJ on this grid is when the solution is completely flat. A flat distribution, when $u_{max}\gg 1$ (highly relativistic) has a maximum temperature $T_{max,rel.}/m_{0}c^{2}\sim u_{max}/2$. However, discretely projecting a flat distribution requires the iterative scheme to drive the input $T$ to infinity. This takes infinitely many iterations, and since we do not want material at the momentum bounds in any case, we settle for a maximum of 20 iterations which allows a $T_{max}/m_{0}c^{2}\sim u_{max}/4.5$ in this limit. Thus, setting an effective maximum temperature.
Likewise, for Figure 3 panel (b), a minimum temperature can be roughly estimated in the relativistic limit. Assuming a tent function in a single cell, $f_{tent}(u)=0.5-|u|/\Delta u$ and integrating between the momentum bounds, ±$\pm\Delta u/2$ gives us a rough minimum temperature estimate by plugging this distribution into the equation 12,
$\frac{T_{min,rel.}}{m_{0}c^{2}}\approx\frac{(\Delta u^{2}+16)\sqrt{4+\Delta u^{2}}-12\text{asinh}(\frac{\Delta u}{2})\Delta u-32}{6\Delta u^{2}}.$ (28)
We plot this estimate of $T_{min,rel.}$ in Figure 3 panel (b).
@@
Figure 3: Example limitations on the MJ temperature supported by the finite velocity grid. Panel (a) shows a non-relativistic limit, while panel (b) considers a relativistic distribution. The rows are ordered in increasing dimensionality of momentum space, from 1D, 2D and 3D. The contours show the number of iterations required for the scheme to converge the moments to an absolute error of $\varepsilon<10^{-12}$. Zero iterations indicate the scheme could not converge the moments within 20 iterations to the required accuracy. As a note for panel (e). The upper left corner, colored white, is simulations which were not run due to the large memory requirements. All panels were run with $n=1$ and $\mathbf{v}_{b}=0$.
A final note for panel (e) when the maximum temperature cannot be resolved. The cause is that for the non-relativistic runs panels (a,c,e) the momentum bounds span approximately 10 thermal velocities. As the dimensionality increased, the relativistic temperatures ($T\sim 1$) at higher dimensionality of the grid, the more material of the distribution is left off and the more iterations are required for the iteration scheme to converge. So, the upper temperature limit which appears on this plot is linked to increasing dimensionality, requiring more iterations to converge. This is similar to the upper temperature limit slope weakening with increased dimensionality.

### v2

@@ 3.2 Robust Algorithm for Computing DG Moments
Numerically, strong gradients in the drift velocity may be super-luminal even if the cell average is less than the speed of light. Constraining the drift velocity to be well-behaved is even harder when trying to compute the DG expanded velocity squared term, $|\mathbf{v}_{b}|^{2}$, in the Lorentz boost factor and the fluid-stationary frame density. To solve this issue, we applied the following algorithm to the relativistic solver to utilize the spatial component of the bulk four-velocity in the moment calculation. The spatial component of the bulk four-velocity is guaranteed to be well behaved as it does not have to be bounded from above, and computing the needed Lorentz boost factor from this quantity is also guaranteed to be positive definite and greater than one. Thus, we obtain an overall robust scheme for computing moments that will not return unphysical gamma factors. To achieve this refactor the steps are now:
@@
    Compute $\mathbf{v}_{b}$ with equation (11) using weak division, equation (21).
@@
      Evaluate $\mathbf{v}_{b}$ at polynomial order 1 Gauss-Lobatto points. (The Gauss-Lobatto points differ from Gauss-Legendre points in that the Gauss-Lobatto points include the cell vertices. Using them ensures physical values at the boundaries and at all points interior points within the cell by using $p=1$ representation.)
@@
    Compute the fluid-stationary frame density, $n$, using weak multiplication: $n=\frac{N}{\Gamma}$.
@@
    $P=nT=\frac{m_{0}}{3}\int f(\mathbf{u})\left(\Gamma_{\mathbf{u}}^{2}\gamma-2\Gamma_{\mathbf{u}}\mathbf{u}_{b}\cdot\mathbf{u}+\frac{(\mathbf{u_{b}}\cdot\mathbf{u})^{2}-1}{\gamma}\right)d^{3}u.$ (23)
@@ 3.3 Maxwell-Jüttner Initialization Routine
Initialization of the MJ distribution begins with the spatially DG-represented quantities $n,\mathbf{v}_{b}$, $T$ and $\Gamma$ computed via the routine in Section 3.2. These quantities are then used to calculate the MJ distribution, equation (2), at the quadrature points. We can then transform the nodal representation onto the modal phase basis, $\phi(\mathbf{x},\mathbf{u})$.
@@
However, directly evaluating the MJ distribution, equation (2), onto the momentum grid has two complications. First, only a limited range of temperatures can be reliably computable because of the exponential behavior of both the distribution’s numerator and the modified Bessel function in the denominator. In particular, the modified Bessel function has significant finite precision errors at low temperatures. Second, in terms of dimension, $d$, the modified Bessel function takes the form $K_{\frac{d+1}{2}}$, which is (potentially) of fractional order. These are notoriously difficult to compute accurately to the desired finite precision.
@@
To avoid the issues of calculating the modified Bessel functions altogether, we replace them with their asymptotic expansions. To leading order, the expansions of $K_{1},K_{3/2},$ and $K_{2}$ are all identical, $K_{\nu}(x)\sim\sqrt{\pi/2x}\exp{(-x)}$ where $\nu=(1,3/2,2)$. They each retain the same leading-order asymptotic behavior as $x\to 0,K_{\nu}(x)\to\infty$ and as $x\to\infty$, $K_{\nu}(x)\to 0$. This means that the projected distribution function is offset in magnitude (the density) only. This unnormalized distribution, expanding $K_{\nu}$ in the Maxwell-Jüttner distribution, equation (2), is written as
@@
$\widetilde{f}^{MJ}(\mathbf{u})=\frac{n}{2(2\pi)^{\frac{d-1}{2}}(m_{0})^{d}\theta^{\frac{d-1}{2}}\sqrt{\pi\theta/2}}\exp\left(\frac{1}{\theta}-\left(\frac{\Gamma(\gamma-\mathbf{v}_{b}\cdot\mathbf{p}/m_{0})}{\theta}\right)\right).$ (24)
[…]
@@
The density moment of equation (24) can be corrected by multiplying the ratio computed via weak division of the correct density moment over the asymptotic density moment to give the properly normalized distribution. Correcting with this density ratio allows us to avoid finite precision errors associated with the modified Bessel functions and simultaneously corrects any errors in the density introduced by the discrete projection.
@@
There are two common cases we consider for this density correction. The first case arises during initial conditions, when we know the input fluid-stationary frame density and thus can simply use that desired density moment to correct the distribution function and avoid errors in both the modified Bessel functions and projections. The second case is for operators such as the BGK collision operator, where the density is a time-dependent quantity. In this case, we compute the moments described in equations (10) and (11) to obtain the fluid-stationary frame density of the evolving distribution function, and then use that density in the correction routine to ensure that the projected MJ distribution has the exact density of the evolving distribution.
@@
In discrete, finite grids, the higher moments of the MJ projection, $\mathbf{v}_{b}$ and $T$, may also deviate significantly from the desired values. To reduce the error between the discretely projected distribution’s moments, we employ an iterative method. Starting with the desired moments, $n_{0}$, $\mathbf{v}_{b,0}$, and $T_{0}$ we create a discrete projection $\tilde{f}$. The moments of this discrete distribution $\tilde{n}$, $\tilde{\mathbf{v}}_{b}$, and $\tilde{T}$, will not necessarily match $n_{0}$, $\mathbf{v}_{b,0}$, and $T_{0}$. We then use Picard iteration to minimize the difference between these sets of moments. For an iteration $k$, and a vector of the moments $\mathbf{M}=\{n,\mathbf{v}_{b},T\}$ we wish to converge, we have an iterations scheme
@@
An alternative algorithm, presented in Section 3.3 of Dzanic et al. (2023), employs a Newton method which, in some cases, would reduce the number of required iterations. This algorithm is extended to the relativistic Maxwell-Jüttner by replacing the distribution function, using the moments outlined in Section 2.1, and recomputing the partials needed by the Jacobian. However, we found that, for modal DG schemes, the Newton method is prohibitively expensive.
For the non-relativistic BGK operator, Dzanic, Witherden and Martinelli Dzanic et al. (2023) report the Newton correction typically takes 1-2 iterations to reach machine precision. Such an approach could also apply to the Maxwell-Jüttner distribution. However, for specifically the modal DG represented quantities, the Jacobian required by this method would be prohibitively expensive to compute, and each cell would need to invert the Jacobian of the size number of basis, times the number of moments, squared. The trade off is that the corrective method provided here will take additional iterations but is simpler to implement and can automatically handle arbitrary equilibrium functions, such as the MJ, Maxwellian, and equilibrium in curved spaces.
@@ 4 Tests
This section explores the accuracy of projecting the MJ distribution function and tests the combination of the projection routine and the moment routine. We address the accuracy of projecting a MJ distribution, examining grid requirements for capturing the distribution within a certain tolerance of error. We demonstrate how the correction routine alters the discrete distribution to match the moments. We also provide analytic estimates for maximum and minimum temperatures representable by the finite momentum grid. Finally, we integrate these concepts into a conservative relativistic BGK operator. All tests in this section were run with the Gkeyll code using polynomial order $p=2$ Serendipity basis. The scripts are publicly available at: https://github.com/ammarhakim/gkyl-paper-inp.
@@ 4.1 Convergence Test
Figure 1: Absolute error between the projected Maxwell-Jüttner distribution’s moments and the desired moments, without the correction routine for higher moments. The moment values here are $n=1.0,v_{b}=0.5$, and $T=1.0$ and the momentum bounds extend from $u_{max}=\pm 160$. At coarser resolutions, the initial non-monotonicity of the velocity error convergence is caused by small differences in the projection of the distribution onto the discrete grid.
@@
From Figure 1, we see that convergence to machine precision without the correction routine of the beam velocity and temperature takes around 10 000 momentum cells with moderate temperature and beam velocity of 1.0 and 0.5, respectively. Convergence indicates we have sufficiently wide momentum bounds to resolve the distribution function for the specified grid size and moments. However, 10 000 grid points is impractical for an actual simulation, especially with higher-dimensional momentum spaces, hence the need for a moment-correction routine. With the correction routine applied to all moments, the same distribution can be represented with only 32 cells, as shown in Figure 4. This reduction and computational savings is one of the central contributions of this paper, making machine precision accurate simulations possible with reasonable domains and with multiple momentum space dimensions.
The requirement for having a large momentum domain of $u_{max}=\pm 160$ in Figure 1 arises from the difference in the MJ distribution’s asymptotic behavior for large values of momentum, $u$. Compared with the Maxwellian distribution, this difference becomes apparent. For a 1D Maxwellian distribution, $f^{Max}\sim\exp(-m_{0}v^{2}/T)$. In contrast, the 1D Maxwell-Jüttner goes as $f^{MJ}\sim\exp(-m_{0}|u|/T)$ when $m_{0}|u|/T\gg 1$. Hence, the Maxwellian approaches zero much faster than the Maxwell-Jüttner on their respective grids. This detail leads to some subtle issues if not minded. To illustrate, the bounds of $u_{max}=\pm 10$ may seem reasonable. However, we are truncating more distribution function than is visually apparent. For example, the lost density fraction for the bounds $u_{max}=\pm 10$ is ($\delta N/N$) $\sim 0.003685$.
@@
Figure 2: MJ plotted with the error between the corrected Maxwell-Jüttner distribution $f_{c}$ and the uncorrected distribution $f_{unc}$. Since the corrected, uncorrected, and theoretic distributions are indistinguishable on the plot, only the corrected distribution is plotted here.
@@ 4.2 Representation Limits of Moments of the MJ Distribution
Figure 3 panel (a) scans the case where the bulk of the distribution is non-relativistic for a 1D momentum space. Contours of the number of iterations required to correct the moments in $\Delta u$ vs temperature $T/m_{0}$ demonstrate that the convergence of temperature is not guaranteed for $T<T_{min,non-rel.}$. Overlaid on the plot, the red line is an estimate of the minimum temperature in the non-relativistic case, which we compute by assuming the distribution is a tent function in a single cell, $f_{tent}(u)=0.5-|u|/\Delta u$, and integrating over the momentum bounds, $\pm\Delta u/2$. Substituting this distribution into equation (12) and taking the non-relativistic limit provides an estimate of the minimum temperature the grid can support. The result is
$\frac{T_{min,non-rel.}}{m_{0}}=\frac{\Delta u^{2}}{24},$ (26)
where $\Delta u$ is the grid spacing in momentum space. Depending on the polynomial order of the solution, the shape of the minimum temperature distribution can vary the constant prefactor. Scanning from warmer to cooler temperatures, the number of iterations required to converge the scheme increases. Eventually, the algorithm hits the maximum number of iterations, and occasionally fails to converge below $T_{min,non-rel.}/m_{0}$. The grid spacing and number of iterations effectively set the minimum grid temperature for the MJ. The trend of minimum temperature holds across 1D, 2D, and 3D momentum spaces.
@@
$\frac{T_{max}}{m_{0}}=\frac{u_{max}\sqrt{1+u_{max}^{2}}-\text{asinh}(u_{max})}{2u_{max}}$ (27)
which arises from computing the temperature of a flat distribution between finite bounds of equation (13) where we have assumed the distribution has no net flow. We also define the inverse hyperbolic trig function as $\text{asinh}(x)=\ln(x+\sqrt{x^{2}+1})$.
In the relativistic limit, Figure 3 panel (b), we can simultaneously show the maximum MJ temperature $T_{max,rel.}/m_{0}$ and the minimum temperature, $T_{min,rel.}/m_{0}$. Scanning $u_{max}$ and $T/m_{0}$, a clear region appears where the temperature is able to converge. Like in the non-relativistic case, the maximum temperature projectable by a MJ on this grid is when the solution is completely flat. A flat distribution, when $u_{max}\gg 1$ (highly relativistic) has a maximum temperature $T_{max,rel.}/m_{0}\sim u_{max}/2$. However, discretely projecting a flat distribution requires the iterative scheme to drive the input $T/m_{0}$ to infinity. This takes infinitely many iterations, and since we do not want material at the momentum bounds in any case, we settle for a maximum of 20 iterations which allows a $T_{max}/m_{0}\sim u_{max}/4.5$ in this limit. Thus, setting an effective maximum temperature.
Likewise, for Figure 3 panel (b), a minimum temperature can be roughly estimated in the relativistic limit. Assuming a tent function in a single cell, $f_{tent}(u)=0.5-|u|/\Delta u$ and integrating between the momentum bounds, $\pm\Delta u/2$ gives us a rough minimum temperature estimate by plugging this distribution into the equation (12),
$\frac{T_{min,rel.}}{m_{0}}\approx\frac{(\Delta u^{2}+16)\sqrt{4+\Delta u^{2}}-12\text{asinh}(\frac{\Delta u}{2})\Delta u-32}{6\Delta u^{2}}.$ (28)
We plot this estimate of $T_{min,rel.}/m_{0}$ as a red line in Figure 3 panel (b).
@@
Figure 3: Example limitations on the MJ temperature supported by the finite momentum grids. The contours show the number of iterations required for the scheme to converge the moments to an absolute error of $\varepsilon<10^{-12}$ at varied temperatures and grid parameters. White regions indicate the correction routine took greater than 20 iterations to converge or was unable to converge. Red lines overlay the temperature-limit estimates from this section. Panels (a), (c), and (e) show the non-relativistic limit, while panels (b), (d), and (f) consider relativistic distributions. The rows are ordered by increasing dimensionality of momentum space, from 1D, 2D and 3D. As a note for panel (e). The upper left corner, colored white, is simulations that were not run due to the large memory requirements. All panels were run with $n=1$ and $\mathbf{v}_{b}=0$.
A final note for panel (e) when the maximum temperature cannot be resolved. The maximum temperature limits for the non-relativistic cases in panels (a), (c), and (e) are due to the fixed width of the momentum bounds at approximately 10 thermal velocities. These extents become insufficient to resolve the distribution as it approaches $T/m_{0}=1$. Particularity as the dimensionality increased, the relativistic temperatures ($T/m_{0}\sim 1$) at higher dimensionality of the grid, the more material of the distribution is left off and the more iterations are required for the iteration scheme to converge. So, the upper temperature limit which appears on this plot is linked to increasing dimensionality, requiring more iterations to converge. This is similar to the upper temperature limit slope weakening with increased dimensionality.
In the white regions of the contours in Figure 3, where the correction routine was unsuccessful, the error between the desired moments and the projected moments begins to rapidly diverge. These areas represent the boundaries of moments resolvable by the grid. Implementing the correction algorithm within the Gkeyll code, we have added checks for convergence between the desired and projected moments at all spatial points and consider the correction routine complete only if all points achieve the desired level of accuracy. If the routine fails, we default to the density-only correction so simulations may continue and have outputs which indicate the failure.
@@
The correction routine developed in this paper could be complemented with additional techniques such as adaptive mesh refinement (AMR) (for a relevant kinetic example, see Kotipalo et al. (2024)), non-uniform momentum-space grids, transforming the kinetic equation into the fluid-stationary frame (see Achterberg & Norman (2018) and Schween & Reville (2024)), or combinations of these techniques. AMR can allow the simulation to flexibly extend its own grid bounds and refine cells if the moments approach the resolvable limits. This would expand the range of the resolvable momentum and temperature moments dynamically, ensuring they successfully converge. Similarly, nonuniform meshes could be employed to coarsely resolve the distribution tail, reducing the truncation error. Lastly, transforming the kinetic equation to evolve in the fluid-stationary frame would eliminate issues with non-symmetric distributions from boosts and eases the grid requirements for challenging configurations such as ultra-relativistic flows. In all of these cases, the correction routine and techniques presented in this paper are still required to maintain conservation laws and enable reasonably compact domains.